dy Solve the initial value problem (t+1) dt (4t2 + 4t) (t+1), y(1) = 7 y(t)
Answer #1

dy dt s 1. tti y 1.E = - (t+1 = y +(4+4 4t) (t+1) ; y(i)=7 4+²+ at er en + (#1) =4+C++) exp [ S- *] ex® [-State ] a exp Elog